OpenStudy (anonymous):

The distance, s(t), in feet traveled by a freely falling object is given by the function s(t) = 16t^2, where t is time in seconds. Use this formula to find the time it would take for an object to fall to the ground from a cliff that is 400 feet high.

OpenStudy (anonymous):

16t^2 = 400 => t^2 = 400/16 => t^2 = 25 => t = 5 seconds.

OpenStudy (anonymous):

wow, that was fast! you are good at math. thank you!:)

OpenStudy (anonymous):

Remember, in this case, even though t can assume + or - 5 in theory, it can't be negative because time can't be negative. So, only positive value is valid.

OpenStudy (anonymous):

Yes, i agree.:)